OpenStudy (anonymous):

Kate borrowed $400 from her friend. She will repay this amount plus 3 percent simple interest after one year. What is the total amount she will pay her friend?

OpenStudy (anonymous):

The formula for simple interest is: \[I = p \times r \times t\] I = Interest p = amount borrowed/principal r = rate (the percent) t = time

OpenStudy (anonymous):

Input the values known in your question into the equation...

OpenStudy (anonymous):

So, it would look like this: \[I = $400 \times 0.03 \times 1\] (Tip: You always change the percent into a decimal. You can do this by moving the decimal place two places to the right.) Solve the equation above to get your answer.

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

Anytime!

OpenStudy (anonymous):

Woops, I noticed an error. You can change a percent to a decimal by moving the decimal point two places to the LEFT.